# Are Graph Neural Networks Optimal Approximation Algorithms?

### v5

@@ C.2 Neural Certification Scheme
Next we establish that for any choice of $\lambda^{*}$ we can compute a lower bound on inequality 108 as follows. Let $F_{\lambda^{*}}(VV^{T})$ be defined as the funciton in the RHS of 108.
@@
Certificate Experiment: We run our certification procedure which we name OptGNN-cert and compare it to the SDP certificate. Note, that mathematically we will always produce a larger (i.e inferior) dual certificate in comparison to the SDP because we are bounding the distance to the SDP optimum with error in the gradients and hessians of the output embeddings of OptGNN. Our advantage is in the speed of the procedure. Without having to go through a primal dual solver, the entire time of producing OptGNN-cert is in the time required to feedforward through OptGNN. In this case we train an OptGNN-Max-Cut with $10$ layers, on $1000$ Erdos-Renyi graphs, with $N=100$ nodes and edge density $p=0.1$. We plot the OptGNN Max-Cut value (an actual integer cut) on the x-axis and in the y-axis we plot the dual certificate value on the same graph where we compare the SDP certificate with the OptGNN-cert. See 3 for the $N=50$ graphs and 2 for the $N=100$ graphs.
Note of course the dual certificate for any technique must be larger than the cut value outputted by OptGNN so the scatter plot must be above the $x=y$ axis of the plot. We see as is mathematically necessary, the OptGNN-cert is not as tight as the SDP certificate but certainly competitive and more importantly it is arrived at dramatically faster. Without any attempt at optimizing the runtime, the OptGNN feedforward and certification takes no more than $0.02$ seconds whereas the SDP takes $0.5$ seconds on $N=100$ node graphs.
@@ Appendix D Experiment details
For Table 3, we display results for forced RB instances drawn from two different distributions. For RB200, we select $N$ uniformly in $[6,15]$ and $K$ uniformly in $[12,21]$. For RB500, we select $N$ uniformly in $[20,34]$ and $K$ uniformly in $[10,29]$.
@@
We also include a greedy baseline, which is the function one_exchange (for Max-Cut) or min_weighted_vertex_cover (for Min-Vertex-Cover) from networkx (Hagberg et al., 2008).
@@
In Table 3 and Table 4, at test time, we use 100 random initial vector assignments instead of just 1, and retain the best score achieved by any hyperplane on any random initial vector assignment. We use 1 random initial vector assignment for validation as in other cases.
@@
Hyperparameters. We ran each experiment on a range of hyperparameters. See Figure 4 for the hyperparameter listing. For all training runs, we used the Adam optimizer Kingma & Ba (2014) with a learning rate of 0.001. We used Laplacian eigenvector Dwivedi et al. (2020) (LE) or random walk Dwivedi et al. (2021) (RW) positional encoding with dimensionality of half the rank, except for rank 32 where we used 8 dimensions. For Max-3-SAT, we set the penalty term $\rho=0.003$.
D.1 Max-3-SAT
ErdősGNN. The ErdősGNN baseline accepts as input a clause-variable bipartite graph of the SAT formula and the node degrees as input attributes. Each variable $x_{i}$ is set to TRUE with probability $p_{i}$. A graph neural network (GatedGCNN) is trained until convergence ($\sim$50k iterations) with randomly generated formulae with clause to variable ratiot in the range $[4,4.3]$. The neural network is trained to produce output probabilities $p_{i}$ by minimizing the expected number of unsatisfied clauses. The exact closed form of the expectation can be found in (Karalias, 2023). At inference time, the probabilities are rounded to binary assignments using the method of conditional expectation.
@@ D.1 Max-3-SAT
D.2 Additional Results
D.2.1 Comparisons with Gurobi and Greedy
@@ D.2.2 Ratio tables
D.3 Model ablation study
@@
Table 7: Performance of various model architectures for selected datasets on Maximum Cut. Here, higher is better. GAT is the Graph Attention network (Veličković et al., 2018), GIN is the Graph Isomorphism Network (Xu et al., 2019), GCNN is the Graph Convolutional Neural Network (Morris et al., 2019), and GatedGCNN is the gated version (Li et al., 2015).
@@
D.4 Effects of hyperparameters on performance
@@ D.5 Out of distribution testing
D.6 Pseudocode for OptGNN training and inference
@@ Appendix E Generalization Analysis
In this section we produce a generalization bound for OptGNN. First we restate our result that a penalized loss approximates the optimum of SDP 1. Our analysis follows from a standard perturbation argument where the key is to bound the lipschitzness of the OptGNN aggregation function. Here we will have to be more precise with the exact polynomial dependence of the lipschitzness of the gradient $\nabla\mathcal{L}_{\rho}$ and the smoothness of the hessian $\nabla^{2}\mathcal{L}_{\rho}$.
@@
We begin by recomputing precisely the number of iterations required for algorithm 2 to approximate the global optimum of SDP 1. Note that this was done in the proof of Theorem C.1 but we do it here with explicit polynomial dependence.
@@
Algorithm 2 computes in $O(\Phi^{4}\epsilon^{-4}\log^{4}(\delta^{-1}))$ iterations a set of vectors $\mathbf{v}\vcentcolon=\{\hat{v}_{s}\}$ for all $s\subseteq\mathcal{S}(P)$ for all $P\in\mathcal{P}$ that satisfy the constraints of SDP 1 to average error $\epsilon$ and approximates the optimum of SDP 1 to error $\epsilon$ with probability $1-\delta$
@@
where $\text{SDP}(\Lambda)$ is the optimum of SDP 1.
@@
To apply the gradient descent lemma of Jin et al. (2017) Theorem F.1 we need a bound on the lipschitzness of the gradient and the smoothness of the hessian of the loss equation 88. By the lipschitz gradient Lemma E.2 we have that the loss is :$\ell\vcentcolon=\text{poly}(B)\rho$ lipschitz, and by the smooth hessian Lemma F.1 we have the loss is $\gamma\vcentcolon=\text{poly(B)}\rho$ smooth. Then we have by Theorem F.1 that perturbed gradient descent in
@@
$=\frac{2\rho}{|\mathcal{P}|}\sqrt{\sum_{w\in\mathcal{F}}\Big{[}\left\|-2\langle v_{w},\delta_{v_{w}}\rangle v_{w}-\left\|\delta_{v_{w}}\right\|^{2}v_{w}-\left\|v_{w}\right\|^{2}\delta_{v_{w}}-2\langle v_{w},\delta_{v_{w}}\rangle\delta_{v_{w}}-\left\|\delta_{v_{w}}\right\|^{2}\delta_{v_{w}}\right\|^{2}\Big{]}}+\frac{2\rho}{|\mathcal{P}|}\|V-\hat{V}\|_{F}$ (186)
@@
$\Pr\left[\left|\frac{1}{\Gamma}\sum_{i\in[\Gamma]}\text{OptGNN}_{(\hat{M},\Lambda_{i})}(V)-\mathbb{E}_{\Lambda\sim\mathcal{D}}\left[\text{OBJ}(\Lambda)\right]\right|\leq\epsilon\right]\geq 1-\delta$
@@ F.1 Miscellaneous Lemmas
Theorem F.1 (perturbed-gd Jin et al. (2017)).
Let $f$ be $\ell$-smooth (that is, it’s gradient is $\ell$-Lipschitz) and have a $\gamma$-Lipschitz Hessian. There exists an absolute constant $c_{max}$ such that for any $\delta\in(0,1),\epsilon\leq\frac{\ell^{2}}{\gamma},\Delta_{f}\geq f(X_{0})-f^{*}$, and constant $c\leq c_{max}$, $PGD(X_{0},\ell,\gamma,\epsilon,c,\delta,\Delta_{f})$ applied to the cost fucntion $f$ outputs a $(\gamma^{2},\epsilon)$ SOSP with probability at least $1-\delta$ in
@@
(Robustness Theorem 4.6 (Raghavendra & Steurer, 2009a) rephrased) Let $\mathbf{v}$ be a set of vectors satisfying the constraints of SDP 1 to additive error $\epsilon$ with objective $OBJ(\mathbf{v})$, then
[…]
@@
The OptGNN of Corollary 2 , which by construction is equivalent to Algorithm 2, outputs a set of embeddings $\mathbf{v}$ such that the rounding of Raghavendra & Steurer (2009a) outputs an integral assignment $\mathcal{V}$ with a Max-k-CSP objective OBJ($\mathcal{V}$) satisfying $OBJ(\mathcal{V})\geq S_{\Lambda}(\text{SDP}(\Lambda)-\epsilon)-\epsilon$ in time $\text{exp}(\text{exp}(\text{poly}(\frac{kq}{\epsilon})))$ which approximately dominates the Unique Games optimal approximation ratio.
@@
The proof follows from the robustness theorem of Raghavendra & Steurer (2009a) which states that any solution to the SDP that satisfies the constraints approximately does not change the objective substantially Theorem F.2. ∎

### v7

@@ C.1 Neural Certification Scheme
Next we establish that for any choice of $\lambda^{*}$ we can compute a lower bound on inequality 92 as follows. Let $F_{\lambda^{*}}(VV^{T})$ be defined as the funciton in the RHS of 92.
@@
Certificate Experiment: We run our certification procedure which we name OptGNN-cert and compare it to the SDP certificate. Note, that mathematically we will always produce a larger (i.e inferior) dual certificate in comparison to the SDP because we are bounding the distance to the SDP optimum with error in the gradients and hessians of the output embeddings of OptGNN. Our advantage is in the speed of the procedure. Without having to go through a primal dual solver, the entire time of producing OptGNN-cert is in the time required to feedforward through OptGNN. In this case we train an OptGNN-Max-Cut with $10$ layers, on $1000$ Erdos-Renyi graphs, with $N=100$ nodes and edge density $p=0.1$. We plot the OptGNN Max-Cut value (an actual integer cut) on the x-axis and in the y-axis we plot the dual certificate value on the same graph where we compare the SDP certificate with the OptGNN-cert. See 2 for the $N=50$ graphs and 1 for the $N=100$ graphs.
Note of course the dual certificate for any technique must be larger than the cut value outputted by OptGNN so the scatter plot must be above the $x=y$ axis of the plot. We see as is mathematically necessary, the OptGNN-cert is not as tight as the SDP certificate but certainly competitive and more importantly it is arrived at dramatically faster. Without any runtime optimizations, the OptGNN feedforward and certification takes no more than $0.02$ seconds whereas the SDP takes $0.5$ seconds on $N=100$ node graphs.
@@ Appendix D Experiment details
For LABEL:fig:forced_rb, we display results for forced RB instances drawn from two different distributions. For RB200, we select $N$ uniformly in $[6,15]$ and $K$ uniformly in $[12,21]$. For RB500, we select $N$ uniformly in $[20,34]$ and $K$ uniformly in $[10,29]$.
@@
We also include a greedy baseline, which is the function one_exchange (for Max-Cut) or min_weighted_vertex_cover (for Min-Vertex-Cover) from networkx (networkx).
@@
In LABEL:fig:forced_rb and LABEL:tab:cut_neural, at test time, we use 100 random initial vector assignments instead of just 1, and retain the best score achieved by any hyperplane on any random initial vector assignment. We use 1 random initial vector assignment for validation as in other cases.
@@
Hyperparameters. We ran each experiment on a range of hyperparameters. See Figure 3 for the hyperparameter listing. For all training runs, we used the Adam optimizer kingma2014adam with a learning rate of 0.001. We used Laplacian eigenvector dwivedi2020benchmarking (LE) or random walk dwivedi2021graph (RW) positional encoding with dimensionality of half the rank, except for rank 32 where we used 8 dimensions. For Max-3-SAT, we set the penalty term $\rho=0.003$.
D.1 Max-Cut
Low-rank SDP.We have included an additional baseline that solves a low-rank SDP using coordinate descent (wang2019low) to the Max-Cut benchmark that originally appears in Any-CSP (tonshoff2022one). We have used the publicly available implementation provided by the authors. For each instance, the rank is automatically selected by the package to be $\lceil\sqrt{2N}\rceil$.
D.2 Max-3-SAT
ErdősGNN. The ErdősGNN baseline accepts as input a clause-variable bipartite graph of the SAT formula and the node degrees as input attributes. Each variable $x_{i}$ is set to TRUE with probability $p_{i}$. A graph neural network (GatedGCNN) is trained until convergence ($\sim$50k iterations) with randomly generated formulae with clause to variable ratiot in the range $[4,4.3]$. The neural network is trained to produce output probabilities $p_{i}$ by minimizing the expected number of unsatisfied clauses. The exact closed form of the expectation can be found in (karalias2023probabilistic). At inference time, the probabilities are rounded to binary assignments using the method of conditional expectation.
Survey propagation. We report the results of survey propagation braunstein2005survey. Typically, algorithms like survey propagation are accompanied by a local search algorithm like WalkSAT. Here we report the results obtained by the algorithm without running any local search on its output.
Low-rank SDP. This is identical to the baseline that we used for Max-3-SAT (see Max-Cut section above).
@@ D.2 Max-3-SAT
D.3 Additional Results
D.3.1 Comparisons with Gurobi and Greedy
@@ D.3.2 Ratio tables
D.4 Model ablation study
@@
Table 5: Performance of various model architectures for selected datasets on Maximum Cut. Here, higher is better. GAT is the Graph Attention network (velivckovic2018graph), GIN is the Graph Isomorphism Network (xu2018how), GCNN is the Graph Convolutional Neural Network (morris2019weisfeiler), and GatedGCNN is the gated version (li2015gated).
@@
D.5 Effects of hyperparameters on performance
@@ D.6 Out of distribution testing
D.7 Pseudocode for OptGNN training and inference
@@ Appendix E Generalization Analysis
In this section we produce a generalization bound for OptGNN. First we restate our result that a penalized loss approximates the optimum of SDP 2. Our analysis follows from a standard perturbation argument where the key is to bound the lipschitzness of the OptGNN aggregation function. Here we will have to be more precise with the exact polynomial dependence of the lipschitzness of the gradient $\nabla\mathcal{L}_{\rho}$ and the smoothness of the hessian $\nabla^{2}\mathcal{L}_{\rho}$.
@@
We begin by recomputing precisely the number of iterations required for algorithm 3 to approximate the global optimum of SDP 2. Note that this was done in the proof of Theorem C.1 but we do it here with explicit polynomial dependence.
@@
Algorithm 3 computes in $O(\Phi^{4}\epsilon^{-4}\log^{4}(\delta^{-1}))$ iterations a set of vectors $\mathbf{v}\vcentcolon=\{\hat{v}_{s}\}$ for all $s\subseteq\mathcal{S}(P)$ for all $P\in\mathcal{P}$ that satisfy the constraints of SDP 2 to average error $\epsilon$ and approximates the optimum of SDP 2 to error $\epsilon$ with probability $1-\delta$
@@
where $\text{SDP}(\Lambda)$ is the optimum of SDP 2.
@@
To apply the gradient descent lemma of jin2017escape Theorem F.1 we need a bound on the lipschitzness of the gradient and the smoothness of the hessian of the loss equation 72. By the lipschitz gradient Lemma E.2 we have that the loss is :$\ell\vcentcolon=\text{poly}(B)\rho$ lipschitz, and by the smooth hessian Lemma F.1 we have the loss is $\gamma\vcentcolon=\text{poly(B)}\rho$ smooth. Then we have by Theorem F.1 that perturbed gradient descent in
@@
$=\frac{2\rho}{|\mathcal{P}|}\sqrt{\sum_{w\in\mathcal{F}}\Big{[}\left\|-2\langle v_{w},\delta_{v_{w}}\rangle v_{w}-\left\|\delta_{v_{w}}\right\|^{2}v_{w}-\left\|v_{w}\right\|^{2}\delta_{v_{w}}-2\langle v_{w},\delta_{v_{w}}\rangle\delta_{v_{w}}-\left\|\delta_{v_{w}}\right\|^{2}\delta_{v_{w}}\right\|^{2}\Big{]}}$ (128)
$+\frac{2\rho}{|\mathcal{P}|}\|V-\hat{V}\|_{F}$ (129)
@@
$\Pr\left[\left|\frac{1}{\Gamma}\sum_{i\in[\Gamma]}\text{OptGNN}_{(\hat{M},\Lambda_{i})}(V)-\mathbb{E}_{\Lambda\sim\mathcal{D}}\left[\text{OptGNN}_{(\hat{M},\Lambda)}(U)\right]\right|\leq\epsilon\right]\geq 1-\delta$
@@ F.1 Miscellaneous Lemmas
Theorem F.1 (perturbed-gd jin2017escape).
Let $f$ be $\ell$-smooth (that is, it’s gradient is $\ell$-Lipschitz) and have a $\gamma$-Lipschitz Hessian. There exists an absolute constant $c_{max}$ such that for any $\delta\in(0,1),\epsilon\leq\frac{\ell^{2}}{\gamma},\Delta_{f}\geq f(X_{0})-f^{*}$, and constant $c\leq c_{max}$, $PGD(X_{0},\ell,\gamma,\epsilon,c,\delta,\Delta_{f})$ applied to the cost function $f$ outputs a $(\gamma^{2},\epsilon)$ SOSP with probability at least $1-\delta$ in
@@
(Robustness Theorem 4.6 (RSanycsp09) rephrased) Let $\mathbf{v}$ be a set of vectors satisfying the constraints of SDP 2 to additive error $\epsilon$ with objective $OBJ(\mathbf{v})$, then
$\text{SDP}(\Lambda)\geq OBJ(\mathbf{v})-\sqrt{\epsilon}poly(kq)$
@@
The OptGNN of Corollary 2 , which by construction is equivalent to Algorithm 3, outputs a set of embeddings $\mathbf{v}$ such that the rounding of RSanycsp09 outputs an integral assignment $\mathcal{V}$ with a Max-k-CSP objective OBJ($\mathcal{V}$) satisfying $OBJ(\mathcal{V})\geq S_{\Lambda}(\text{SDP}(\Lambda)-\epsilon)-\epsilon$ in time $\text{exp}(\text{exp}(\text{poly}(\frac{kq}{\epsilon})))$ which approximately dominates the Unique Games optimal approximation ratio.
@@
The proof follows from the robustness theorem of RSanycsp09 which states that any solution to the SDP that satisfies the constraints approximately does not change the objective substantially Theorem F.2. ∎
Appendix G Definitions
In this section we introduce precise definitions for message passing, message passing GNN, and SDP’s. First we begin with the SDP.
Definition (Standard Form SDP).
An SDP instance $\Lambda$ is comprised of objective matrix $C\in\mathbb{R}^{N\times N}$ and constraint matrices $\{A_{i}\}_{i\in\mathcal{F}}$ and constants $\{b_{i}\}_{i\in\mathcal{F}}$ over a constraint set $\mathcal{F}$
Minimize: $\langle C,X\rangle$ (158)
$X\succeq 0$ (160)
For any standard form SDP, there is an equivalent vector form SDP with an identical optimum.
Definition (Vector Form SDP).
Any standard form SDP $\Lambda$ is equivalent to the following vector form SDP.
Minimize: $\langle C,V^{T}V\rangle$ (161)
Subject to: $\langle A_{i},V^{T}V\rangle=b_{i}\quad\forall i\in[\mathcal{F}]$ (162)
$V=[v_{1},v_{2},...,v_{N}]\in\mathbb{R}^{N\times N}$ (163)
For the SDP’s corresponding to CSP’s, there is a natural graph upon which the CSP instance is defined. Next we define what it means for a message passing algorithm on a graph to solve a SDP.
Definition (Message Passing).
A $T$ iteration message passing algorithm denoted MP is a uniform circuit family that takes as input a graph $G=(V,E)$ and initial embeddings $U=\{u^{0}_{i}\}_{i\in[N]}\in\mathbb{R}^{r\times N}$ and outputs $\text{MP}(G,U^{0})\in\mathbb{R}^{r\times N}$. The evaluation of MP involves the use of a uniform circuit family defined for each iteration $\text{UPDATE}=\{\text{UPDATE}_{j}\}_{j\in[T]}$. At iteration $\ell\in[T]$, for each node $i\in V$, the function $\text{UPDATE}_{\ell}:\mathbb{R}^{r\times(|N(i)|+1)}\rightarrow\mathbb{R}^{r}$ takes the embeddings of nodes $i$ and its neighbors at iteration $\ell-1$, denoted $\{u^{\ell-1}_{j}\}_{j\in N(i)\cup i}\in\mathbb{R}^{r}$, and outputs embedding $u^{\ell}_{i}\in\mathbb{R}^{r}$.
We additionally require some mild restrictions on the UPDATE function to be polynomially smooth in its inputs and computable in polynomial time. That is for any $\|U\|\leq B$ in the $B$ norm ball we have.
$\text{UPDATE}_{r}(U-U^{\prime})\leq poly(B,N)\|U-U^{\prime}\|$ (164)
And the $\text{UPDATE}_{\ell}$ circuit is polynomial time computable $\mathrm{poly}(|N(i)|,r)$.
We impose some mild restrictions on the form of the message passing algorithm to capture algorithms that could reasonably be executed by a Message Passing GNN. In practice the UPDATE circuit for OptGNN is a smooth cubic polynomial similar to linear attention and therefore of practical value. Next we define Message Passing GNN.
Definition (Message Passing GNN).
A $T$ layer Message Passing GNN is a uniform circuit family that takes as input a graph $G=(V,E)$, a set of parameters $M\vcentcolon=\{M^{1},M^{2},...,M^{T}\}$, and initial embeddings $U^{0}=\{u^{0}_{i}\}_{i\in V}\in\mathbb{R}^{r\times N}$ and outputs $\text{GNN}(G,M,U^{0})\in\mathbb{R}^{r\times N}$. The GNN circuit is evaluated as follows. For each node $i\in V$, each layer $\ell\in[T]$, there is a uniform circuit family $\text{AGG}_{\ell}:\mathbb{R}^{|M^{\ell}|}\times\mathbb{R}^{r\times|N(i)|}\rightarrow\mathbb{R}^{r}$ that takes as input a set of parameters $M^{\ell}\in M$ and a set of embeddings of node $i$ and its neighbors at layer $\ell-1$ denoted $\{u^{\ell-1}_{j}\}_{j\in N(i)\cup i}$, and outputs an embedding $u^{\ell}_{i}$. This update equation is represented as follows.
$u^{\ell}_{i}\vcentcolon=\text{AGG}(M^{\ell},U^{\ell})=\text{UPD}(u^{\ell-1}_{i},\text{MSG}(\{u^{\ell-1}_{j}\}_{j\in N(i)}))$ (165)
For some functions UPD and MSG parameterized by $M^{\ell}$. To capture meaningful models of GNNs, we require $\text{AGG}(M,U)$ to be polynomially lipschitz and computable in polynomial time. That is, for weights $\hat{M}$ and $M^{\prime}$ and for inputs $\hat{U}$ and $U^{\prime}$ all in the $B$ norm ball, we require
$\|\text{AGG}(\hat{M},\hat{U})-\text{AGG}(M^{\prime},U^{\prime})\|\leq\mathrm{poly}(B,N)\left(\|\hat{M}-M^{\prime}\|+\|\hat{U}-U^{\prime}\|\right)$ (166)
and we require $\text{AGG}(\hat{M},\hat{U})$ to be computable in $poly(|N(i)|,r)$ time.
In practice the AGG circuit for OptGNN is a smooth cubic polynomial similar to linear attention and therefore of practical value. In almost the same way, we define a message passing algorithm see Definition Definition.
Definition (Output of GNN solves SDP).
We say a GNN $\epsilon$-approximately solves an SDP with optimum OPT if its output embeddings $U^{\ell}=\{u_{i}\}_{i\in[N]}$ approximately optimize the vector form SDP objective $\text{SDP}(\{u_{i}\}_{i\in[N]})\geq\text{OPT}-\epsilon$ and approximately satisfy the vector form constraints $\left|\langle A_{i},V^{T}V\rangle-b_{i}\right|\leq\epsilon$